OpenStudy (anonymous):

Which of the following equations best represents the graph of f(p)? y = f(p) I think it's this one f(y) = p f(x) = p x = f(p)

OpenStudy (anonymous):

@phi

OpenStudy (anonymous):

@ganeshie8

OpenStudy (anonymous):

You picked the right one

OpenStudy (anonymous):

I have a couple more,

OpenStudy (anonymous):

If f(x) = x - 4, what does f(3) represent? The value of (x - 4) when x = 3 The value of x when (x - 4) = 3 The value of 3f(x) when x = 3 The value of f(x) when (x - 4) = 3

OpenStudy (anonymous):

which one do you think it is

OpenStudy (anonymous):

@ganeshie8

OpenStudy (anonymous):

f(3) means that x=3

OpenStudy (anonymous):

you just plug in 3 where ever x is